Plasma Quest update ...

Li-Wen Chang lwchang at
Fri Aug 22 18:38:19 PDT 2008

I have ran two wafers (1 dummy + 1 carrier wafer with sample) at 100C
chuck temp.
Still the clamp down sensor did not work properly.
For the dummy wafer everything looked fine (3sccm backside He for
10Torr, Bias -61.5V for oxygen3 and -48.3V for my rcp, which is
However for my 2nd carrier wafer the bias was very low ~
-15.6V.(Backside He was 3sccm) The RF wattage was stable, though

Again I've noticed wafer popping out the pings despite cooling the
chuck temp to 50C before unloading.


On Fri, Aug 22, 2008 at 3:44 PM, John Shott <shott at> wrote:
> Plasma Quest community:
> This afternoon Elmer and I spent some time trying to resolve some of the outstanding issues on the Plasma Quest.
> In particular, we changed the clamping springs to a weaker set that, we believe, still clamp hard limit helium blow by, but won't clamp so hard that we break wafers.
> We also worked to treat the O-ring in an oxygen plasma in hopes that this will tend to limit wafer sticking to the O-ring and the subsequent drop off the pins.
> We have put everything back together and a releasing things for the weekend with the following cautions:
> 1. Please keep an eye on your wafers during load and unload to watch for wafers falling off pins and other things that might result in wafer breakage.
> 2. Please try to note helium flow to maintain 10 Torr backside pressure and the bias observed during your process.  While we think that the helium flow is behaving pretty well, bias may still be somewhat low.
> 3. If you run into problems and get a trapped wafer, send me email.  I should be available many times during the weekend to extract wafers, etc.
> Note: if the oxygen-plasma-treated O-ring doesn't reduce sticking problems, I've also got a couple of spare O-rings that I can try Elmer's special bead-blasting-to-reduce-sticking treatment.
> So best of luck this weekend .... and let me know if you encounter problems.
> John

More information about the pquest mailing list